# On the exponents of distribution of primes and smooth numbers

### v1

@@ 1. Introduction
$\lambda_{q}=\sum_{q=q_{1}q_{2}q_{3}}\alpha_{q_{1}}\beta_{q_{2}}\gamma_{q_{3}}.$
@@
The analogous equidistribution problem for smooth numbers [21, 11] concerns the quantities
@@
where $P^{+}(n)$ denotes the largest prime factor of $n$; the key setting is $y\leq x^{1/C}$, with $C$ large. In this context, Granville [23, Theorem 2] proved a suitable analogue of the Bombieri–Vinogradov theorem, achieving the exponent of distribution $\tfrac{1}{2}-\varepsilon$ (see also [46, 47, 20, 24]). Relying on a triple convolution estimate of Bombieri–Friedlander–Iwaniec [4, Theorem 4], Fouvry–Tenenbaum [21] raised the exponent to $\tfrac{3}{5}-\varepsilon$, with an upper bound of $x(\log x)^{-A}$ as in 1.1. Drappeau later [11] strengthened the bound back to $\Psi(x,y)(\log x)^{-A}$, with the same exponent of $\tfrac{3}{5}-\varepsilon$. We remark that all of these results use absolute values (i.e., arbitrary $1$-bounded weights $\lambda_{q}$), which is possible beyond the square-root barrier due to the flexible factorization properties of smooth numbers.
@@
One can extend Theorem 1.5 to smooth-supported multiplicative functions, using our triple convolution estimate (Proposition 6.3) in the work of Drappeau–Granville–Shao [12].
@@
#$\#\{n\leq x:P^{+}(n),P^{+}(n+1)\leq y\}\ll_{\varepsilon}\varrho(u)^{1+5/8-\varepsilon},$
@@ 2.2. Reaching the exceptional spectrum
$\sum_{r\sim Q^{4}/x^{2}}\sum_{f}^{\Gamma_{0}(r)}\sqrt{x}^{\theta_{f}}\left|\sum_{m\sim Q^{2}/x}\rho_{f}(m)\sum_{n\sim Q^{4}/x^{2}}a_{n,r}\overline{\rho}_{f}(n)\right|\stackrel{{\scriptstyle?}}{{<}}\frac{Q^{3}}{x},$ (2.9)
@@
where we really need a power-saving in the final bound. Plugging in Selberg’s bound $\theta\leq 1/2$ (which was the state of the art in [8]), one reaches the exponent of distribution $\tfrac{5}{8}-o(1)$ from the works of Maynard [37] and Drappeau [11]. Using the celebrated bound $\theta\leq\tfrac{7}{32}$ of Kim–Sarnak [30, Appendix 2], one reaches the exponent $\tfrac{66}{107}-o(1)$ from the works of Lichtman [32] and the author [39]. Conditionally on Selberg’s conjecture that $\theta=0$, the resulting exponent is $\tfrac{5}{8}-o(1)$.
@@ 4. Primes with triply-well-factorable weights
where we omitted a factor of $Q^{-3\theta/2}$ in the second term of the expansion. For this to be acceptable in 4.9 (i.e., $\ll_{\varepsilon,a}x^{-7\varepsilon}N^{2}R^{2}S^{3}/d$), we need the following restrictions:
@@ 5. Primes with linear sieve weights
$\mathbf{D}_{r}^{-}(D)$ :=$:=\left\{(P_{1},\ldots,P_{r}):P_{1}\geq\cdots\geq P_{r}\geq 1,\,P_{1}\cdots P_{j-1}P_{j}^{3}\leq D\text{ for even }j\leq r,\,D_{1}^{2}\leq D\right\},$
@@ 5.1. The upper-bound linear sieve weights
for all $20\leq k<r/$ 2 and $1\leq k\leq r$. We aim to place the primes $p_{1},\ldots,p_{r}$ into $d_{1},d_{2},d_{3}$ such that the bounds in 5.3 hold.
@@
Case 3.1: If $d_{2}>x^{1/5-5\delta}$ and $d_{3}\leq x^{1/3+2\delta}/N$, we set $D_{1}:=Nx^{-\delta}$, $D_{2}:=d_{2}$, and $D_{3}:=x^{3/5-3\delta}/(Nd_{2})$, which have $D_{i}\geq d_{i}$ (in light of 5.6) and $D_{1}D_{2}D_{3}\geq D$. We then run a greedy process to place the remaining primes $p_{k}$ with $k\geq j+1$ into either $d_{1}$ or $d_{3}$, while preserving the inequalities $d_{i}\leq D_{i}$. This works because at step $k$, before placing $p_{k}$, we have
@@
Case 3.2: Otherwise, we have $d_{2}\leq x^{1/5-5\delta}$ and $d_{3}>x^{1/3+2\delta}/N$. We then set $D_{1}:=Nx^{-\delta}$, $D_{2}:=x^{3/5-3\delta}/(Nd_{3})$, and $D_{3}:=d_{3}$, which have $D_{i}\geq d_{i}$ (in light of 5.6) and $D_{1}D_{2}D_{3}\geq D$. We run a similar greedy process to place the remaining primes $p_{k}$ with $k\geq j+1$ into either $d_{1}$ or $d_{2}$, while preserving the bounds $d_{i}\leq D_{i}$. This works because at step $k$, before placing $p_{k}$, we have
@@ 5.2. The linear sieve weights with special factors
where $\psi(x,y):=x\mathbbm{1}_{x\geq y+4\delta}$ and
$w(t_{1},t_{2},t_{3})=\psi\left(\min\left\{\frac{5-3t_{3}}{8},1-2t_{2}\right\},\frac{1+t_{1}}{2}\right).$
@@
Moreover, if $t_{1}\leq\tfrac{1}{4}$ and $r\geq 3$, then it suffices that $\vartheta_{0}\leq\vartheta(t_{1},t_{2},t_{3})-4\delta$, as in 5.8.
@@
Let $\delta,\vartheta,u,v,N$ be as in Lemma 5.7, $r\geq 3$, $1\leq D\leq x^{\vartheta-2\delta}$, and $d\in\mathcal{D}^{\textnormal{well}}(D)$. Write $d=p_{1}\cdots p_{r}$ where $p_{1}\geq\cdots\geq p_{r}$ are primes. Assume that $p_{3}\leq x^{u-v}$, and ($p_{1}\leq x^{v}$, $p_{2}^{2}\leq x^{1-\vartheta-2\delta}$) or ($p_{2}\leq x^{v}$, $p_{1}^{2}\leq x^{1-\vartheta-2\delta}$). Then there exists a factorization $d=d_{1}d_{2}d_{3}$ into positive integers satisfying 5.3.
@@
Let $d=p_{1}\cdots p_{r}$ where $p_{1}\geq\cdots\geq p_{r}$ are primes with $p_{i}=x^{t_{i}}$. Consider the set $I$ of all exponents $\vartheta_{0}\in[0,1)$ such that if $d\in\mathcal{D}^{\textnormal{well}}(x^{\vartheta_{0}})$, then there exists a factorization $d=d_{1}d_{2}d_{3}$ into positive integers obeying 5.3. Since $\mathcal{D}^{\textnormal{well}}(D)\subset\mathcal{D}^{\textnormal{well}}(D^{\prime})$ when $D\leq D^{\prime}$, $I$ is an interval of the form $[0,\vartheta_{\max})$; our goal is to show that $\vartheta(t_{1})-2\delta\in I$ as in 5.7, and moreover, if $t_{1}\leq\tfrac{1}{4}$ and $r\geq 3$, then $\vartheta(t_{1},t_{2},t_{3})-4\delta\in I$ as in 5.8.
Applying Proposition 5.9 for some $\vartheta\in[\tfrac{1}{2},\tfrac{5}{8}]$, we deduce that $\vartheta-2\delta\in I$ provided $t_{1}\leq u$, and that one of the following holds:
@@
Recalling the definitions of $\vartheta(t)$ and $u,v$ from 5.9 and 5.7, a short computation gives
$t\in[u,v]=\left[2\vartheta-1,\frac{2-2\vartheta}{3}\right]\qquad\iff\qquad\vartheta\leq\vartheta(t)=\min\left(\frac{1+t}{2},\frac{2-3t}{2}\right).$
Using $t=t_{1}$ in $(i)$, which automatically satisfies $t_{1}\leq u$, we deduce that
$\vartheta(t_{1})-2\delta\in I,$
as required. Now assume that $t_{1}\leq\tfrac{1}{4}$ and $r\geq 3$; then we have $t_{1}\leq\tfrac{1}{4}\leq(2-2\vartheta)/3=u$ for any $\vartheta\leq\tfrac{5}{8}$, so the condition $t_{1}\leq u$ holds automatically. Thus points $(i)$-$(v)$ above imply, respectively:
  • $(i)$.
    $\max\{\vartheta(t_{1}),\vartheta(t_{2}),\vartheta(t_{1}+t_{2}),\vartheta(t_{1}+t_{2}+t_{3})\}-2\delta\in I$;
  • $(ii)$.
    $\vartheta(t_{1}+t_{3})-2\delta\in I$ if the range $[t_{1}+2t_{2}+t_{3}+2\delta,\vartheta(t_{1}+t_{3})]$ is nonempty;
  • $(iii)$.
    $\vartheta(t_{2}+t_{3})-2\delta\in I$ if the range $[2t_{1}+t_{2}+t_{3}+2\delta,\vartheta(t_{2}+t_{3})]$ is nonempty;
  • $(iv)$.
    $\min\left\{\tfrac{5-3t_{3}}{8},1-2t_{2}-2\delta\right\}-2\delta\in I$ if $[\tfrac{1+t_{1}}{2},\min\left\{\tfrac{5-3t_{3}}{8},1-2t_{2}-2\delta\right\}-2\delta]$ is nonempty;
  • $(v)$.
    $\min\left\{\tfrac{5-3t_{3}}{8},1-2t_{1}-2\delta\right\}-2\delta\in I$ if $[\tfrac{1+t_{2}}{2},\min\left\{\tfrac{5-3t_{3}}{8},1-2t_{1}-2\delta\right\}-2\delta]$ is nonempty.
Putting these together, using the cutoff function $\psi(x,y)=x\mathbbm{1}_{x\geq y+4\delta}$ to encode the nonemptiness of the required ranges above, and recalling 5.8, we conclude that
$\vartheta(t_{1},t_{2},t_{3})-4\delta\in I,$
as required. ∎
Finally, we prove Corollary 1.4 on the count of twin primes up to $x$.

### v2

@@ 1. Introduction
$\lambda_{q}=\sum_{q_{1}q_{2}q_{3}=q}\alpha_{q_{1}}\beta_{q_{2}}\gamma_{q_{3}}.$
@@
The analogous equidistribution problem for smooth (friable) numbers [21, 11] concerns the quantities
@@
where $P^{+}(n)$ denotes the largest prime factor of $n$; here $y\leq x^{1/C}$, with $C$ large. In this context, Granville [23, Theorem 2] proved a suitable analogue of the Bombieri–Vinogradov theorem, achieving the exponent of distribution $\tfrac{1}{2}-\varepsilon$ (see also [46, 47, 20, 24]). Relying on a triple convolution estimate of Bombieri–Friedlander–Iwaniec [4, Theorem 4], Fouvry–Tenenbaum [21] raised the exponent to $\tfrac{3}{5}-\varepsilon$, with an upper bound of $x(\log x)^{-A}$ as in 1.1. Drappeau later [11] strengthened the bound back to $\Psi(x,y)(\log x)^{-A}$, with the same exponent of $\tfrac{3}{5}-\varepsilon$. We remark that all of these results use absolute values (i.e., arbitrary $1$-bounded weights $\lambda_{q}$), which is possible beyond the square-root barrier due to the flexible factorization properties of smooth numbers.
@@
Following Drappeau–Granville–Shao [12], one can deduce a similar result for smooth-supported multiplicative functions in arithmetic progressions, using a slight extension of our triple convolution estimate (Proposition 6.3).
@@
#$\#\{n\leq x:P^{+}(n),P^{+}(n+1)\leq y\}\ll_{\varepsilon}x\varrho(u)^{1+5/8-\varepsilon},$
@@
1.1. Acknowledgements
We thank James Maynard, Sary Drappeau, Jori Merikoski, Lasse Grimmelt, and Jared Duker Lichtman for many helpful comments. The author is sponsored by the EPSRC Scholarship at University of Oxford.
@@ 2.2. Reaching the exceptional spectrum
$\sum_{r\sim Q^{4}/x^{2}}\sum_{f}^{\Gamma_{0}(r)}\sqrt{x}^{\theta_{f}}\left|\sum_{m\sim Q^{2}/x}\rho_{f}(m)\sum_{n\sim Q^{4}/x^{2}}a_{n,r}\overline{\rho}_{f}(n)\right|<\frac{Q^{3}}{x},$ (2.9)
@@
where we really need a power-saving in the final bound. Plugging in Selberg’s bound $\theta\leq 1/2$ (which was the state of the art in [8]), one reaches the exponent of distribution $\tfrac{3}{5}-o(1)$ from the works of Maynard [37] and Drappeau [11]. Using the celebrated bound $\theta\leq\tfrac{7}{32}$ of Kim–Sarnak [30, Appendix 2], one reaches the exponent $\tfrac{66}{107}-o(1)$ from the works of Lichtman [32] and the author [39]. Conditionally on Selberg’s conjecture that $\theta=0$, the resulting exponent is $\tfrac{5}{8}-o(1)$.
@@ 4. Primes with triply-well-factorable weights
where we omitted a factor of $Q^{-3\theta/2}$ in the first term of the expansion. For this to be acceptable in 4.9 (i.e., $\ll_{\varepsilon,a}x^{-7\varepsilon}N^{2}R^{2}S^{3}/d$), we need the following restrictions:
@@ 5. Primes with linear sieve weights
$\mathbf{D}_{r}^{-}(D)$ :=$:=\left\{(P_{1},\ldots,P_{r}):P_{1}\geq\cdots\geq P_{r}\geq 1,\,P_{1}\cdots P_{j-1}P_{j}^{3}\leq D\text{ for even }j\leq r,\,P_{1}^{2}\leq D\right\},$
@@ 5.1. The upper-bound linear sieve weights
for all $20\leq j<r/$ 2 and $1\leq k\leq r$. We aim to place the primes $p_{1},\ldots,p_{r}$ into $d_{1},d_{2},d_{3}$ such that the bounds in 5.3 hold.
@@
Case 3.1: One has $d_{2}>x^{1/5-5\delta}$ and $d_{3}\leq x^{1/3+2\delta}/N$. In this case, we set $D_{1}:=Nx^{-\delta}$, $D_{2}:=d_{2}$, and $D_{3}:=x^{3/5-3\delta}/(Nd_{2})$, which have $D_{i}\geq d_{i}$ (in light of 5.6) and $D_{1}D_{2}D_{3}\geq D$. We then run a greedy process to place the remaining primes $p_{k}$ with $k\geq j+1$ into either $d_{1}$ or $d_{3}$, while preserving the inequalities $d_{i}\leq D_{i}$. This works because at step $k$, before placing $p_{k}$, we have
@@
Case 3.2: One has $d_{2}\leq x^{1/5-5\delta}$ and $d_{3}>x^{1/3+2\delta}/N$. Then we set $D_{1}:=Nx^{-\delta}$, $D_{2}:=x^{3/5-3\delta}/(Nd_{3})$, and $D_{3}:=d_{3}$, which have $D_{i}\geq d_{i}$ (in light of 5.6) and $D_{1}D_{2}D_{3}\geq D$. We run a similar greedy process to place the remaining primes $p_{k}$ with $k\geq j+1$ into either $d_{1}$ or $d_{2}$, while preserving the bounds $d_{i}\leq D_{i}$. This works because at step $k$, before placing $p_{k}$, we have
@@ 5.2. The linear sieve weights with special factors
where $\psi(x,y):=x\mathbbm{1}_{x\geq y+2\delta}$ and
$w(t_{1},t_{2},t_{3})=\psi\left(\min\left\{\frac{5-3t_{3}}{8},1-2t_{2}-2\delta\right\},\frac{1+t_{1}}{2}\right).$
@@
Moreover, if $t_{1}\leq\tfrac{1}{4}$ and $r\geq 3$, then it suffices that $\vartheta_{0}\leq\vartheta(t_{1},t_{2},t_{3})-2\delta$, as in 5.8.
@@
Let $\delta,\vartheta,u,v,N$ be as in Lemma 5.7, $r\geq 3$, $1\leq D\leq x^{\vartheta-2\delta}$, and $d\in\mathcal{D}^{\textnormal{well}}(D)$. Write $d=p_{1}\cdots p_{r}$ where $p_{1}\geq\cdots\geq p_{r}$ are primes. Suppose that $p_{3}\leq x^{u-v}$. Also, assume that ($p_{1}\leq x^{v}$, $p_{2}^{2}\leq x^{1-\vartheta-2\delta}$) or ($p_{2}\leq x^{v}$, $p_{1}^{2}\leq x^{1-\vartheta-2\delta}$). Then there exists a factorization $d=d_{1}d_{2}d_{3}$ into positive integers satisfying 5.3.
@@
Let $d=p_{1}\cdots p_{r}$ where $p_{1}\geq\cdots\geq p_{r}$ are primes with $p_{i}=x^{t_{i}}$. We want to show that $d$ has a factorization obeying 5.3, under one of the following assumptions:
  • $(a)$.
    $d\in\mathcal{D}^{\textnormal{well}}(x^{\vartheta(t_{1})-2\delta})$ where $\vartheta(t_{1})$ is as in 5.7, or
  • $(b)$.
    $t_{1}\leq\tfrac{1}{4}$, $r\geq 3$, and $d\in\mathcal{D}^{\textnormal{well}}(x^{\vartheta(t_{1},t_{2},t_{3})-2\delta})$, where $\vartheta(t_{1},t_{2},t_{3})$ is as in 5.8.
Applying Proposition 5.9 for some $\vartheta\in[\tfrac{1}{2},\tfrac{5}{8}]$ and letting $u=u(\vartheta)$, $v=v(\vartheta)$ be as in 5.9, we deduce that $d$ has a factorization obeying 5.3 provided that $d\in\mathcal{D}^{\textnormal{well}}(x^{\vartheta-2\delta})$, $t_{1}\leq u$, and that one of the following holds:
@@
Note that from 5.9, 5.7 and a short computation, we have the equivalence
$t\in[v,u]=\left[2\vartheta-1,\frac{2-2\vartheta}{3}\right]\qquad\iff\qquad\vartheta\leq\vartheta(t)=\min\left(\frac{1+t}{2},\frac{2-3t}{2}\right).$
Now suppose assumption $(a)$ holds. Then we can use $\vartheta=\vartheta(t_{1})$, which implies $t_{1}\in[v,u]$. So $d\in\mathcal{D}^{\textnormal{well}}(x^{\vartheta-2\delta})$, $t_{1}\leq u$, and $(i)$ holds for $t=t_{1}$, and thus $d$ has a factorization as required.
Next, suppose assumption $(b)$ holds. Then we can use $\vartheta=\vartheta(t_{1},t_{2},t_{3})$, which also lies in $[\tfrac{1}{2},\tfrac{5}{8}]$. Moreover, we have $t_{1}\leq\tfrac{1}{4}\leq\tfrac{2-2\vartheta}{3}=u$. So $d\in\mathcal{D}^{\textnormal{well}}(x^{\vartheta-2\delta})$, $t_{1}\leq u$, and it suffices to verify one of conditions $(i)$-$(v)$ above; we split into cases based on the maximum from 5.8:
  • If $\vartheta\in\{\vartheta(t_{1}),\vartheta(t_{2}),\vartheta(t_{1}+t_{2}),\vartheta(t_{1}+t_{2}+t_{3})\}$, then $(i)$ holds;
  • If $\vartheta=\psi(\vartheta(t_{1}+t_{3}),t_{1}+2t_{2}+t_{3})$, so $\vartheta=\vartheta(t_{1}+t_{3})$ and $t_{1}+2t_{2}+t_{3}+2\delta\leq\vartheta(t_{1}+t_{3})$, then $(ii)$ holds;
  • If $\vartheta=\psi(\vartheta(t_{2}+t_{3}),2t_{1}+t_{2}+t_{3})$, so $\vartheta=\vartheta(t_{2}+t_{3})$ and $2t_{1}+t_{2}+t_{3}+2\delta\leq\vartheta(t_{2}+t_{3})$, then $(iii)$ holds;
  • If $\vartheta=w(t_{1},t_{2},t_{3})$, so $\vartheta=\min\left\{\tfrac{5-3t_{3}}{8},1-2t_{2}-2\delta\right\}$ and $\tfrac{1+t_{1}}{2}+2\delta\leq\vartheta$, then $(iv)$ holds (noting that $u-v=\frac{5-8\vartheta}{3}$);
  • If $\vartheta=w(t_{2},t_{1},t_{3})$, so $\vartheta=\min\left\{\tfrac{5-3t_{3}}{8},1-2t_{1}-2\delta\right\}$ and $\tfrac{1+t_{2}}{2}+2\delta\leq\vartheta$, then $(v)$ holds.
This completes our proof. ∎